Everdeen, INC has a 100 day operating cycle. If it's average age of inventory is 35 days.

How long is its average collection period?

If its average payment period is 30 days what is its cash conversion cycle?
Answer #1

operating cycle=average age of inventory+average collection period

average collection period=(100-35)=65 days

cash conversion cycle=average age of inventory+average collection period-average payment period

=35+65-30=70 days